# Wavespeed selection of travelling wave solutions of a two-component reaction-diffusion model of cell invasion

### v1

@@ 4.2 Inner problem and matching
In the inner region, $U$ and $W$ are small, and $V=O(1)$. Let $U=\gamma^{-1}\hat{U}$, $W=\gamma^{-1}\hat{W}$. To leading order in this inner region, the system (9) is
$\hat{U}^{\prime}=\hat{W},\quad V^{\prime}=\frac{1}{2}\hat{U}V,\quad\hat{W}^{\prime}=\frac{1}{1-V}\left(\frac{1}{2}\hat{U}V\hat{W}-2\hat{W}-\hat{U}(1-V)\right).$ (22)
Considering solutions for which $V_{\infty}=V_{s}$, the relevant inner boundary condition is that the trajectory approaches the given $V=V_{\infty}$ along the more negative eigenvector, that is,
@@
$\hat{W}^{\prime}\sim-2\hat{W}-\hat{U},\qquad\hat{U}^{\prime}=\hat{W},\qquad z\to-\infty.$
@@
[…]
$\log(A_{I})=\lim_{z\to-\infty}\frac{\hat{U}}{\hat{U}+\hat{W}}+\log|\hat{U}+\hat{W}|.$
@@
$\left.A_{I}\right|_{V_{\infty}=0.25}\approx 0.515,\qquad\left.A_{I}\right|_{V_{\infty}=0.5}\approx 1.485,\qquad\left.A_{I}\right|_{V_{\infty}=0.75}\approx 2.943.$ (26)
In order for the inner solution (25) to match to the intermediate region (21), the sinusoidal term in the intermediate solution must approach zero, which happens when its argument is in the neighbourhood of $\pi$. We define a new spatial variable $\xi$ by $z-z_{0}=\delta^{-1/2}\pi+\xi$. In matching we then require $z_{I}=z_{0}+\delta^{-1/2}\pi$, and
$-A_{0}\mathrm{e}^{-\pi\delta^{-1/2}}\xi\mathrm{e}^{-\xi}\sim-A_{I}\gamma^{-1}\xi\mathrm{e}^{-\xi},$
@@
Figure 6: Comparison between the numerically determined travelling wave speeds and the asymptotic approximation for large $\gamma$. (a) Comparison between the wavespeed correction $\delta=2-c$ and numerically determined values and the leading order term in (27) show convergence to the leading order universal logarithmic behaviour that is independent of $V_{\infty}$. (b) Comparison between the next term in the asymptotic approximation (27), in which the effect of $V_{\infty}$ plays a role.
Finally, we note that in the intermediate region in which $U$ behaves as (21), both $U$ and $V$ are small. This region therefore corresponds to the interstitial region that can be observed in Fig. 5. The asymptotic analysis predicts that the width of this region is $O(\delta^{-1/2})=O(\log\gamma)$, hence the slow growth of this region in the large $\gamma$ limit.
@@ 5 Discussion
In this article we have determined the properties of the three dimensional phase space that select the wave speed of the system (5) for compactly supported or exponentially decaying initial conditions. Although the phase-space properties of the three-dimensional system in travelling wave coordinates (9) are more complicated than the two-dimensional Fisher–KPP system, the general principle of wave speed selection is the same. For an initial condition of given far field exponential decay, a PDE solution will evolve toward the travelling wave solution which has the same decay rate, if such a nonnegative solution exists. Otherwise, the solution will tend to the travelling wave solution which has the largest decay rate. Depending on the value of $\gamma$ and $v_{0}$, the solution with fastest decay rate will either correspond to the wave speed $c$ such that $v_{0}=V_{c}$, or such that $v_{0}=V_{s}$, if that point exists. An important point regarding the determination of $V_{s}$ is, although it is a property of the phase space, it is not a local property in the neighbourhood of the $V$-axis; rather, it is related to the manner in which the two-dimensional unstable manifold of $(U,V,W)=(1,0,0)$ intersects the $V$-axis. For this reason, the existence or location of $V_{s}$ cannot be determined by a local expansion, so there is unlikely to be a general closed-form expression for the selected wave speed.
@@
We have also performed the asymptotic analysis of travelling waves in the large $\gamma$ limit. In carrying out this limit it is useful to think of the system as a generalisation of the Fisher–KPP system in which the reaction term is perturbed, leading to a logarithmically small correction to the wave speed from the critical Fisher–KPP wave speed of $2$. The leading order correction in this case is universal, so is likely to be the correction observed for a large class of systems that look like (5) with variations in diffusivity and reaction terms; for example, the system considered by Colson et al. [7], in which the resident cell population effects the diffusivity but not the reaction, is likely to have the same leading order asymptotic behaviour.
In this work we have not considered the degenerate-diffusion case where $v_{0}=1$ (or $V_{\infty}=1$ in calculating travelling wave solutions). In this case, as discussed in El-Hachem et al. [11] and in the closely related problem in Colson et al. [7], the appropriate trajectories end at the point $(U,V,W)=(0,1,0)$, which is a removable singularity in the phase space (as it is the intersection of the $V$-axis and the plane $V=1$), rather than a fixed point. Travelling waves therefore do not behave exponentially near this point. A local expansion will give the permitted behaviour near this point; one method of recovering this is to desingularise the system (9) and perform a centre manifold analysis, similar to that considered in Gallay and Mascia [14], for example. It is unclear if, in this case, the selected wave speed $c$ (corresponding to a compactly supported initial condition) will be selected by a mechanism similar to that which determines the point $V_{s}$: that is, out of all trajectories over $c$ that end at $(0,1,0)$, there will be a unique trajectory such that the decay is maximum. We note that the PDE simulations in El-Hachem et al. [11] suggest that the wavespeed at $v_{0}=1$ is also the one approached if one takes the limit as $v_{0}\to 1$ from below, which indicates that the wave speed selection mechanism should be related.

### v2

@@ 4.2 Inner problem and matching
In the inner region, $V=O(1)$ is required to satisfy the far field condition $V\to V_{\infty}$ as $z\to\infty$. From (11b), this is only possible if $U$ (and thus $W$) is small, in particular of order $\gamma^{-1}$.
Formally, let $U=\gamma^{-1}\hat{U}$, $W=\gamma^{-1}\hat{W}$. To leading order (for large $\gamma$), the system (11) is
$\frac{\mathrm{d}\hat{U}}{\mathrm{d}z}=\hat{W},\quad\frac{\mathrm{d}V}{\mathrm{d}z}=\frac{1}{2}\hat{U}V,\quad\frac{\mathrm{d}\hat{W}}{\mathrm{d}z}=\frac{1}{1-V}\left(\frac{1}{2}\hat{U}V\hat{W}-2\hat{W}-\hat{U}(1-V)\right).$ (26)
Considering solutions for which $V_{\infty}=V_{s}$, the relevant boundary condition in this region is that the trajectory approaches the given $V=V_{\infty}$ along the more negative eigenvector, that is,
@@
$\frac{\mathrm{d}\hat{W}}{\mathrm{d}z}\sim-2\hat{W}-\hat{U},\qquad\frac{\mathrm{d}\hat{U}}{\mathrm{d}z}=\hat{W},\qquad z\to-\infty.$
@@
As with the analysis of the outer problem, the constant $A_{I}$ depends in a nontrivial way on the entire solution to (26), and thus on $V_{\infty}$. The translation constant $z_{I}$ may be thought of as a reference location in the inner region, where $\hat{U}$ takes a specific value (we omit the specific numerical calculation here, as it also depends on $V_{\infty}$).
With the sign convention in (29), $A_{I}$ will be positive. By integrating (26) numerically in the negative $z$ direction, from an initial condition (27), we may numerically compute $A_{I}$ independently of $z_{I}$ in a similar way as was done in the outer problem (22):
$\log(A_{I})=\lim_{z\to-\infty}\left[\frac{\hat{U}}{\hat{U}+\hat{W}}+\log\left\lvert\hat{U}+\hat{W}\right\rvert\right$].
@@
$\left.A_{I}\right\rvert_{V_{\infty}=0.25}\approx 0.515,\qquad\left.A_{I}\right\rvert_{V_{\infty}=0.5}\approx 1.485,\qquad\left.A_{I}\right\rvert_{V_{\infty}=0.75}\approx 2.943.$ (30)
For completeness, in Fig. 8 we plot numerically calculated values of $A_{I}$ over a range of $V_{\infty}$.
Figure 8: Numerically calculated prefactor $A_{I}$ in the far field behaviour (29) of the inner problem (26), as dependent on $V_{\infty}$. The constant $A_{I}$ is required in order to perform the matching with the outer problem, and ultimately determine the first correction to the wave speed $c$ as a function of parameter $\gamma$ and $V_{\infty}$.
We now carry out the matching between the intermediate and inner solutions. In order for the inner solution (29) to match to the intermediate solution (25), the sinusoidal term in the intermediate solution must approach zero, which first happens when its argument is in the neighbourhood of $\pi$. We define a new spatial variable $\xi$ by $z-z_{0}=\delta^{-1/2}\pi+\xi$. Then expanding the intermediate solution in small $\delta$ with $\xi=O(1)$ we obtain
$U_{M}\sim A_{0}\delta^{-1/2}\mathrm{e}^{-\delta^{-1/2}\pi-\xi}\sin(\pi+\delta^{1/2}\xi)\sim-A_{0}\mathrm{e}^{-\delta^{-1/2}\pi}\xi\mathrm{e}^{-\xi}.$
To match this expression with (29) we require $\xi=z-z_{I}$, so that
$z_{I}-z_{0}\sim\delta^{-1/2}\pi.$ (31)
Equating the prefactors between $U_{M}$ and the inner solution $\gamma^{-1}\hat{U}$, we find
$A_{0}\mathrm{e}^{-\pi\delta^{-1/2}}\sim-A_{I}\gamma^{-1},$
@@
Figure 9: Comparison between the numerically determined travelling wave speeds and the asymptotic approximation for large $\gamma$. (a) Comparison between the numerically determined wavespeed correction $\delta=2-c$ (solid lines) and the leading order term in (32) (dashed line) show convergence to the leading order universal logarithmic behaviour that is independent of $V_{\infty}$. (b) The difference between the numerically determined wave speed correction and the leading order approximation (solid lines), compared with the next term in the asymptotic approximation (32) (dashed lines), in which the effect of $V_{\infty}$ plays a role. These results are calculated up to $\gamma=10^{11}$, at which point $\log(\gamma)]^{-2}\approx 0.0015$.
Finally, we note that in the intermediate region in which $U$ behaves as (25), both $U$ and $V$ are small. This region therefore corresponds to the interstitial region that can be observed in Fig. 5. Combining (31) with the asymptotic expression for $\delta$ (32) predicts that the width of this region to leading order is
$z_{I}-z_{0}\sim\log(\gamma),$ (33)
hence the slow growth of this region in the large $\gamma$ limit.
@@ 5 Discussion
In this article we have determined the properties of the three dimensional phase space that select the wave speed $c$ of the system (8) for compactly supported or exponentially decaying initial conditions. Although the phase-space properties of the three-dimensional system in travelling wave coordinates (11) are more complicated than the two-dimensional Fisher–KPP system, the general principle of wave speed selection is the same. For an initial condition of given far field exponential decay, a PDE solution will evolve toward the travelling wave solution which has the same decay rate, if such a nonnegative solution exists, leading to the formula (14). Otherwise, the solution will tend to the travelling wave solution which has the largest decay rate. Depending on the values of $\gamma$ and $v_{0}$, the solution with fastest decay rate will either correspond to the wave speed $c$ such that $v_{0}=V_{c}$, where $V_{c}$ is the threshold between monotonic and oscillatory decay at the front (17), or such that $v_{0}=V_{s}$, where $v_{s}$ separates exponential decay from above and below, if that point exists. An important point regarding the determination of $V_{s}$ is, although it is a property of the phase space, it is not a local property in the neighbourhood of the $V$-axis; rather, it is related to the manner in which the two-dimensional unstable manifold of $(U,V,W)=(1,0,0)$ intersects the $V$-axis. For this reason, the existence or location of $V_{s}$ cannot be determined by a local expansion, so there is unlikely to be a general closed-form expression for the selected wave speed.
@@
We have also performed the asymptotic analysis of travelling waves in the large $\gamma$ limit. In this limit we have focused on solutions selected by $V=V_{s}$, as particular relevant for compactly supported initial conditions, although this is the appropriate value for $\gamma\to\infty$ for any initial condition with sufficiently large decay such that the wave speed $c<2$ (that is, the constraint (15) is not satisfied). In carrying out the $\gamma\to\infty$ limit it is useful to think of the system (11) as a generalisation of the Fisher–KPP system in which the reaction term is perturbed, leading to a logarithmically small correction to the wave speed from the critical Fisher–KPP wave speed of $2$. The leading order correction in this case is universal, so is likely to be the correction observed for a large class of systems that look like (8) with variations in diffusivity and reaction terms; for example, the system considered by Colson et al. [7], in which the resident cell population effects the diffusivity but not the reaction, is likely to have the same leading order asymptotic behaviour.
More generally, the universality of the perturbed Fisher–KPP asymptotic structure means that it plays a role in the asymptotic structure of more complex models, such as the Gatenby and Gawlinski [15] model (5). The logarithmic dependence of gap width on cell degradation rate certainly appears universal. Specific details such as prefactors can depend on the distinguished limits considered, however. The leading-order asymptotic approximation for the interstitial gap width ($z_{+}$) found in Gatenby and Gawlinski [15] and Fasano et al. [13], and established rigorously in [10], is
$z_{+}\sim\frac{1}{\sqrt{\gamma_{2}}}\log\alpha,\qquad d\to 0,$
where we have used the nondimensional variables in the scaling (5). Importantly, the limit in which this holds is vanishing diffusivity ratio $d\to 0$, where $\gamma_{2}$ is taken to be $O(d)$; the prefactor of $1/\sqrt{\gamma_{2}}$ then arises from the balance between the reaction and diffusion terms for the acid concentration in (5). In contrast, the two-component models such as that considered in our study are derived by assuming $\gamma_{2}\gg 1$, so that the invading cell population and acid concentration are effectively coupled. The logarithmic dependence of interstitial gap separation on degradation rate is maintained, however.
From a biological perspective, the results of the above asymptotic analysis imply that when the interstitial gap is appreciable, there is only very weak connection between the interstitial gap width and the wave speed. The invading wave speed is essentially set by the invasion dynamics of the cancer cells in the absence of any resistance by crowding of resident cells (that is, approximately the Fisher–KPP wave speed). In this case, the wave speed is related to the invading cell parameters only (that is, the invading cell diffusivity and proliferation rates), while the interstitial gap width provides only order-of-magnitude information on the effect of invading cells on the environment, including both the excess acidity, and the effect of that acidity on resident cells and extracellular matrix. Thus, while the presence of an interstitial gap may indicate more aggressive tumour growth, there is little difference between different width gaps if they are present.
In this work we have not considered the degenerate-diffusion case where the initial resident cell population is $v_{0}=1$ (or $V_{\infty}=1$ in calculating travelling wave solutions), at which point the diffusivity of invading cells vanishes. In phase space such a trajectory would have to either approach $(U,V,W)=(0,1,0)$ as $z\to\infty$ (if solutions remain smooth), or a non-smooth condition where $(U,V)\to(0,1)$ as $z$ tends to some finite value, with $W$ either approaching a constant or becoming unbounded. For the first such case, it can be shown by substituting an algebraic expansion into (11) that there is a solution that behaves like
$U\sim\frac{2c^{2}}{\gamma}z^{-2},\qquad V\sim 1-2cz^{-1},\qquad W\sim-\frac{4c^{2}}{\gamma}z^{-3},\qquad z\to\infty.$
However, such a solution decays algebraically as $z\to\infty$, which is not the behaviour seen in the numerical simulations; instead, one would expect degenerate diffusion to result in a solution that is non-smooth at a moving front [11]. We have not been able to ascertain a consistent expansion of the system (11) with such behaviour, which we leave for future exploration.
Data availability
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Declarations
Conflict of interest
The authors declare that they have no conflicts of interest.
Ethical statement
This study did not involve any experiments with human participants or animals.